OpenStudy (anonymous):

if v1=(2,-5) and v2=(4,-3) then the angle between the two vectors is_____ degrees. (Round your answer to two decimal places)

OpenStudy (anonymous):

@Hero can you help me?

OpenStudy (irishboy123):

dot product them

OpenStudy (anonymous):

would it be -23

OpenStudy (irishboy123):

that's part of the process but where did you get minus sign?

OpenStudy (anonymous):

well when you add-8 and -15, you get -23, or did I do it wrong?

OpenStudy (irishboy123):

-8 and - 15?? same again, where are the signs coming from ?!

OpenStudy (irishboy123):

2*4 + (-5)*(-3)

OpenStudy (anonymous):

what I did was (-2)(4)+(5)(-3) and that is how I -23

OpenStudy (irishboy123):

\(\vec a \bullet \vec b = <a_{x}, a_{y}>\bullet <b_{x},b_{y}> = a_{x}*b_{x}+a_{y}*b_{y}\)

OpenStudy (irishboy123):

you have \(\vec v_1=(2,-5)\) and \( \vec v_2=(4,-3)\)

OpenStudy (anonymous):

actually they are (-2,5) and v2=(4,-3),

OpenStudy (irishboy123):

can you just check that again because you posted differently above...

OpenStudy (anonymous):

oh never mind, I am going stupid, sorry about that, I got mixed up!

OpenStudy (irishboy123):

so is the opening post right, then?

OpenStudy (irishboy123):

or is it (-2,5) and v2=(4,-3) ?

OpenStudy (anonymous):

yeah the opening post is that right one, soory again!

OpenStudy (irishboy123):

so the dot prod is 23, right?!?! next, the dot product is also this: \(\vec a \bullet \vec b = |\vec a | \ | \vec b| \ cos \theta\), right? so you need \( |\vec a |\) and \(| \vec b|\)....

OpenStudy (irishboy123):

\(|\vec a| = \sqrt {a_x^2 + a_y ^2}\)

OpenStudy (irishboy123):

do that for v1, and v2

OpenStudy (irishboy123):

and then \(\large \theta = cos^{-1}( \frac {23}{|\vec v_1| \ \vec v_2|})\) gotta dash...

OpenStudy (anonymous):

This is hard i am getting really confused!

OpenStudy (anonymous):

Wait for v1 and v2 in the bottom of the equation, what do i sub them in for?

OpenStudy (irishboy123):

i showed you above how to calc \(|\vec a|\), which is the magnitude of \(\vec a\) you do that for v1 and v2 then put them into the equation i'll do v1 \( |\vec v_1| = \sqrt { 2^2 + (-5)^2} = \sqrt{4 + 25} = \sqrt {29}\) can you do the same for v2?